OpenStudy (mathgenus123):

plz help asap i will medal and fan Which equation is X =5 a solution? Show all work for both. * A. 2x - 50 = -8x OR B. 10x/5 = -10

OpenStudy (anonymous):

A. 2(5)-50=-8(5) 10-50=-40 -40=-40 B. \[(10\times5)/5=-10\] 50/5=-10 \[10\neq-10\]